OpenStudy (anonymous):

One-hundred students were allowed to re-take an exam for their math course. The probability distribution shows how studying for the latest exam affected their grade when compared with the first time they took the exam. What is the probability that a student who studied for the exam saw an increase in their exam grade? Round to the nearest thousandth. Exam Grades Studied Did Not Study Totals Raise in Grade 0.52 0.06 0.58 No Raise in Grade 0.05 0.37 0.42 Totals 0.57 0.43 1

OpenStudy (anonymous):

you need to format the data better so we can help you

OpenStudy (anonymous):

right now it makes no sense

OpenStudy (anonymous):

take a picture or draw it

OpenStudy (anonymous):

I know that's why I am to but I go to online school and that is how it came up on the screen

OpenStudy (anonymous):

can somebody please help me figure this out please??

OpenStudy (anonymous):

sorry no, just take a screenshot. use the tool on windows, or on mac press command-shift-4 and drag over what you want to click

OpenStudy (anonymous):

like that

OpenStudy (anonymous):

how do I take a screenshot on windows???

OpenStudy (anonymous):

programs >> accessories >> snipping tool

OpenStudy (anonymous):

from what i remember, i don't know windows 8 if you have that

OpenStudy (anonymous):

I use prt scrn > paint > paste > select the problem > crop > save > attach

OpenStudy (anonymous):

here watch this video: http://www.capture-screenshot.org/snipping-tool/ or read the short instructions

ganeshie8 (ganeshie8):

\(\large \begin{array}{|c|c||c|} \hline \text{Exam Grades}&\text{Studied}&\text{Did Not Study}&\text{Totals}\\ \hline \text{Raise in Grade}&0.52&0.06&0.58\\ \hline \text{No Raise in Grade}&0.05&0.37&0.42\\ \hline \text{Totals}&0.57&0.43&1\\ \hline \end{array} \)

ganeshie8 (ganeshie8):

something like this ?

OpenStudy (anonymous):

that's what I thought it would be

ganeshie8 (ganeshie8):

Okay, and your question is : `What is the probability that a student who studied for the exam saw an increase in their exam grade? `

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

So only look at the "\(\text{Studied}\)" column

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I took the screen shot

OpenStudy (anonymous):

i chose B 0.897 and it was marked wrong

ganeshie8 (ganeshie8):

Probability for \(\text{"Studied" AND "Raise in Grade"} = 0.52 \) Probability for \(\text{"Studied" } = 0.57 \) \(\large \implies \) Probability for \(\text{Raise in Grade}\) given that he \(\text{Studied} = \dfrac{0.52}{0.57}\)

OpenStudy (anonymous):

oh wow your teacher sucks, what a terribly formatted question :( i'm sorry, you should mention that the homework is not formatted well

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

i did and they said there was nothing they could do

OpenStudy (anonymous):

C .912

OpenStudy (anonymous):

dang. well yeah @ganeshie8 sounds right, conditional probability problem

ganeshie8 (ganeshie8):

yeah bayee's thm.. in a convoluted format lol

OpenStudy (anonymous):

thank you guys for taking time and helping me you don't even know how much i appreciate it :D thank you!

OpenStudy (anonymous):

I did nothing :) maybe taught you about screenshots lol

OpenStudy (anonymous):

lol hey you helped me learn something though

OpenStudy (wolf1728):

I found an answer of .912

OpenStudy (anonymous):

yeah i had help already but thank you so much for helping still

OpenStudy (wolf1728):

Okay wesniki23

OpenStudy (anonymous):

Hey just a tip to elaborate on how @ganeshie8 derived 0.52/0.57: So, essentially you have two things relevant to the question. Students who studied, and students who raised their grade. Let's calls students who studied: group A Let's call students who raised their grade: group B Baye's Theorm states that: \[P(B|A) = \frac{P(AB)}{P(A)}\] Meaning, the probability of a student who raised their grade, given that they studied ... is equal to... the probability of students who studided AND raised their grade... divided by.... the probability of students who studied this is derived from: P(AB) = P(A)P(B}A) which is just the basic multiplication rule to find the intersection of A and B (students who studied and raised their grade), you multiply the probability of those who studied by those who raised their grade, given that they studied. algebraically manipulating this equation to solve for P(B}A) is what @ganeshie8 did

OpenStudy (anonymous):

The P(AB) = P(A)P(B|A) is the same as P(AB) = P(A)*P(B) as long as they are independent events. The P(B|A) accounts for dependence